# NeuroAMP: A Novel End-to-end General Purpose Deep Neural Amplifier for Personalized Hearing Aids

### v1

@@ Abstract
The prevalence of hearing aids is increasing. However, optimizing the amplification processes of hearing aids remains challenging due to the complexity of integrating multiple modular components in traditional methods. To address this challenge, we present NeuroAMP, a novel deep neural network designed for end-to-end, personalized amplification in hearing aids. NeuroAMP leverages both spectral features and the listener’s audiogram as inputs, and we investigate four architectures: Convolutional Neural Network (CNN), Long Short-Term Memory (LSTM), Convolutional Recurrent Neural Network (CRNN), and Transformer. We also introduce Denoising NeuroAMP, an extension that integrates noise reduction along with amplification capabilities for improved performance in real-world scenarios. To enhance generalization, a comprehensive data augmentation strategy was employed during training on diverse speech (TIMIT and TMHINT) and music (Cadenza Challenge MUSIC) datasets. Evaluation using the Hearing Aid Speech Perception Index (HASPI), Hearing Aid Speech Quality Index (HASQI), and Hearing Aid Audio Quality Index (HAAQI) demonstrates that the Transformer architecture within NeuroAMP achieves the best performance, with SRCC scores of 0.9927 (HASQI) and 0.9905 (HASPI) on TIMIT, and 0.9738 (HAAQI) on the Cadenza Challenge MUSIC dataset. Notably, our data augmentation strategy maintains high performance on unseen datasets (e.g., VCTK, MUSDB18-HQ). Furthermore, Denoising NeuroAMP outperforms both the conventional NAL-R+WDRC approach and a two-stage baseline on the VoiceBank+DEMAND dataset, achieving a 10% improvement in both HASPI (0.90) and HASQI (0.59) scores. These results highlight the potential of NeuroAMP and Denoising NeuroAMP to deliver notable improvements in personalized hearing aid amplification.
@@
This research introduces NeuroAMP, a novel deep learning-based hearing aid amplifier that replaces complex multi-stage signal processing with streamlined, end-to-end personalized amplification. We also develop Denoising NeuroAMP, which integrates noise reduction along with the amplification process. Validated on diverse datasets, these models outperform conventional and two-stage methods on standardized metrics such as HASPI and HASQI scores. Furthermore, we believe NeuroAMP’s data-driven approach can contribute to advancements in hearing aid technology, particularly by enhancing user-centered hearing assistance and potentially supporting more effective communication for individuals with hearing loss. Future work will focus on real-time implementation and integrating user feedback to further refine personalization.
@@
hearing aids, deep neural network, end-to-end amplification, NeuroAMP, Denoising
@@ 1 Introduction
Hearing loss, a pervasive global health issue, continuously impacts the lives of hundreds of millions, leading to social isolation [1], depression [2], diminished quality of life [3], and even cognitive decline [4]. The World Health Organization (WHO) highlights that approximately 430 million people worldwide currently face problems related to hearing loss [5, 6]. Despite the availability of hearing aids (HAs), which can mitigate these adverse effects, only 16% of hearing-loss individuals in the United States consistently use them [7, 8]. This low adoption rate suggests a notable gap between the potential of hearing aids and their real-world utilization, emphasizing the importance of developing more effective and user-friendly solutions.
@@
Non-use of hearing aids by people with hearing loss can be attributed to a variety of factors, including hearing aids being ineffective in noisy environments, suboptimal sound quality, insufficient benefit, and incompatibility with the individual’s specific type of hearing loss [9]. To address these challenges, various amplifier strategies for hearing aids have been developed [10, 11, 12, 13, 14]. These strategies primarily involve two key components: the prescription fitting formulas and compression techniques to determine the desired insertion gain, as shown in Fig.  1. The prescription formula aims to enhance the audibility of audio based on an individual’s audiogram [10, 11]. Well-known prescription fitting formulas include NAL-NL1 [12], NAL-NL2 [15], NAL-R [10], and DSL m[i/o] [16].
@@
A common problem for people with hearing loss is a reduction in the auditory dynamic range, known as loudness recruitment. In this case, weak sounds may become inaudible, while loud sounds can be uncomfortably intense. Non-linear prescription formulas like NAL-NL1, NAL-NL2, and DSL m[i/o] play a crucial role in addressing the reduced auditory dynamic range through compression. These advanced formulas, however, are primarily used in commercial hearing aids and require licensing, which may limit broader accessibility. In this study, we utilize the widely adopted NAL-R prescription formula and incorporate it with Wide Dynamic Range Compression (WDRC) [14] to introduce non-linearity, mimicking the behavior of more advanced systems like NAL-NL1. This approach aligns with the established practice in the open-source OpenMHA platform [17]. Moreover, while conventional amplifiers demonstrate notable performance, they often face challenges in adapting to dynamic acoustic environments.
Figure 1: Overview of a typical hearing-aid system, illustrating the modular components involved in signal processing, including microphone input, amplification, compression, and speaker output.
@@
These limitations in adaptability and optimization have led to the exploration of alternative approaches. One such promising approach is deep learning, which offers the potential to learn complex patterns from large datasets and create more adaptable systems. Recently, with the availability of huge amounts of training data and their corresponding labels, deep learning models have achieved notable performance in several speech processing applications. Additionally, the integration of deep learning models for speech processing in hearing aid devices has demonstrated robust performance, for example, in speech enhancement [18, 19] and speech assessment tasks [20, 21, 22].
@@
Due to the advancement of deep learning, in this paper, we propose a deep learning-based neural amplifier model called NeuroAMP, designed to provide personalized amplification in an end-to-end manner. Our initial goal is to demonstrate that NeuroAMP can achieve comparable performance to established methods, while also providing the potential for future improvements through its inherent adaptability. Unlike conventional amplifiers (as indicated by the red dashed box in Fig. 1) that combine prescription fitting formulas and compression techniques to amplify audio inputs, NeuroAMP uses a neural network to perform amplification in an end-to-end manner. We hypothesized that the non-linearity of the neural network would allow NeuroAMP to capture a greater variety of acoustic information and more effectively adjust gain and compression based on the user’s specific degree of hearing loss at different frequencies. The development of NeuroAMP involved exploring various neural network architectures, including convolutional neural networks (CNN), long short-term memory (LSTM), convolutional recurrent neural networks (CRNN), and Transformers. The training objective of NeuroAMP is to minimize the loss between the estimated amplified audio and the corresponding ground-truth audio. To ensure robust performance across diverse conditions, we also apply a data-augmentation strategy using datasets containing multiple languages and music. Experimental results confirm the superiority of the Transformer model over other models, achieving the best performance in almost all metrics (with Spearman’s Rank Correlation Coefficient (SRCC) scores of 0.9927 for Hearing Aid Speech Quality Index (HASQI) and 0.9905 for Hearing Aid Speech Perception Index (HASPI) on the TIMIT dataset, and 0.9738 for HAAQI on the Cadenza dataset). Additionally, we also verify the effectiveness of data augmentation in maintaining good performance on unseen datasets (with SRCC scores of 0.9810 for HASQI and 0.9914 for HASPI on the VCTK dataset, and 0.9892 for HAAQI on the MUSDDB18-HQ dataset). Furthermore, we demonstrate that NeuroAMP is able to mimic conventional amplification methods and provide personalized amplification in an end-to-end manner.
To demonstrate its integration potential, we further introduce Denoising NeuroAMP, which combines NeuroAMP with a speech enhancement module, showcasing direct integration between amplification and enhancement. The development of Denoising NeuroAMP involves combining amplification and noise reduction in one module. Specifically, the model takes noisy audio input and outputs enhanced-amplified audio, and the training objective is to minimize the loss between the predicted enhanced-amplified output from the noisy input and the corresponding ground-truth clean audio. Experimental results confirm that Denoising NeuroAMP can achieve a HASPI score of 0.90 and a HASQI score of 0.59, compared to 0.89 and 0.55 for NAL-R+WDRC, and 0.85 and 0.30 for the two-stage Denoising>NAL-R+WDRC baseline, which further confirms the advantages of Denoising NeuroAMP as an integrated amplification and denoising approach.
@@
The remainder of this paper is organized as follows. First, we review related work in Section 2. Second, we introduce the proposed methods in Section 3. In Section 4, we describe the experimental setup, report the experimental results, and discuss our findings. Finally, we conclude our work in Section 5.
@@ 2.1 Prescription Fitting Formulas
Prescription fitting formulas aim to amplify sound based on the user’s degree of hearing loss, making speech signals audible across different frequencies while maintaining comfort. These formulas can be broadly categorized into linear and non-linear approaches. Linear amplification, exemplified by methods like NAL-R and NAL-RP [23], provides a constant gain regardless of input volume and is generally preferred for individuals with flatter hearing loss profiles, larger dynamic ranges, and less variability in dynamic range across frequencies. It is also well-suited for those with less varied lifestyles and environments [13]. NAL-R, developed by the Australian National Acoustic Laboratories, is commonly used for mild to moderate hearing loss, while NAL-RP is tailored for severe hearing loss. In contrast, non-linear amplification, such as NAL-NL1 [12], NAL-NL2 [15], and DSL m[i/o] [16], adjusts the gain dynamically based on the input volume, frequency, and the user’s specific hearing loss characteristics. This makes it more suitable for individuals with sloping hearing loss, reduced dynamic ranges, significant variations in dynamic range across frequencies, and more diverse lifestyles and listening environments. NAL-NL1 aims to equalize and normalize the relative volume levels of individual frequencies while optimizing the speech intelligibility for a given volume [12]. NAL-NL2, an improved version of NAL-NL1, considers additional factors in calculating gains, such as gender, age, and whether the native language is tonal, which may affect output volume. NAL-NL2 is also the most widely used fitting formula at present. In contrast, DSL m[i/o] is designed to prevent uncomfortable loudness while using hearing aids and to maximize the audibility of essential messages in conversations. Furthermore, these non-linear formulas utilize compression techniques to manage the reduced auditory dynamic range commonly experienced by individuals with sensorineural hearing loss.
@@ 2.2 Hearing Aid Compression
People with sensorineural hearing loss tend to have a smaller auditory dynamic range. For these individuals, soft sounds are inaudible, while intense sounds are perceived as loudly as they are by a normal-hearing ear. To optimize the use of the residual dynamic range and restore normal loudness perception, most modern hearing aids adopt compression. Wide Dynamic Range Compression (WDRC) [14] is a widely adopted compression strategy designed to prevent sounds from being either too loud or too soft. The use of WDRC in hearing aids provides several benefits, such as ensuring consistent audibility of speech signals at safe and comfortable hearing levels [14]. Moreover, WDRC involves multiple steps of audio processing, which are [24]:
@@ 3.1 NeuroAMP Architecture
The processing workflow of NeuroAMP involves taking an audio signal and the listener’s audiogram as inputs, processing them through a neural network, and producing an amplified output signal. The detailed architecture of the NeuroAMP model is illustrated in Fig. 2. Specifically, given the input signal $y$, the STFT is processed to extract $y$ into its time-frequency spectral features ($\textbf{Y}=[\textbf{y}_{1},\dots,\textbf{y}_{t},\dots,\textbf{y}_{T}]$). Simultaneously, the audiogram $z$ is transformed into a vector representation via a dense layer and then replicated into a sequence of length $T$ ($\textbf{Z}=[\textbf{z}_{1},\dots,\textbf{z}_{t},\dots,\textbf{z}_{T}]$). The two input streams are concatenated along the y-axis feature dimension and then passed to the subsequent layers of the model. In our setup, we plan to concatenate two sets of features in the latent space, leveraging its ability to capture semantic feature similarity and perform feature fusion. This approach has been shown to be effective for building personalized deep-learning models [25]. The feedforward process of NeuroAMP is defined as follows:
$\begin{array}[]{c}\mathbf{Y}=\text{STFT}(y)\\ \mathbf{Z}=\text{Dense}(z)\\ \mathbf{Concat}=[\mathbf{Y}\mid\mathbf{Z}]\\ \mathbf{\hat{Y}}=\text{NeuroAMP}(\mathbf{Concat}).\\ \end{array}$ (1)
@@
In this work, for the core NN module in NeuroAMP, we select four neural network architectures, including the Convolutional Neural Network (CNN), Long Short-Term Memory (LSTM), Convolutional Recurrent Neural Network (CRNN), and Transformer.
@@ 3.1.1 Convolutional Neural Network (CNN)
The Convolutional Neural Network (CNN) is integrated into the NeuroAMP model due to its ability to extract hierarchical features from spatial or spectral data [26]. CNNs are particularly adept at recognizing patterns and features in audio spectrograms from complex audio signals [27, 28]. In the NeuroAMP model, the CNN is used to process spectral features extracted by the STFT, with the goal of capturing detailed representations of the frequency components of the audio signal. The operation process of the CNN in the NeuroAMP model can be described by the following equations:
@@
$\vdots$
$\mathbf{\hat{Y}}$ $=\text{Dense}(\text{Flatten}(\mathbf{h}_{L-1}))$
where $\mathbf{W}_{l}$ and $\mathbf{b}_{l}$ are the convolutional weights and biases of the $l$-th layer, respectively, and $L$ represents the number of convolutional layers. This architecture ensures a comprehensive feature extraction process, allowing the NeuroAMP model to more effectively distinguish between auditory elements.
@@ 3.1.2 Long Short-Term Memory (LSTM)
LSTM is chosen for its proficiency in modeling temporal dependencies in sequential data [29, 30]. In the context of hearing aids, preserving the temporal structure of the audio signal is essential to maintaining speech intelligibility and naturalness. LSTM is effective at capturing long-range dependencies and temporal patterns [31], which is crucial for processing continuous audio signals and adapting to varying acoustic environments. The process of the LSTM in the NeuroAMP model is described as follows:
$\mathbf{h}_{t},\mathbf{c}_{t}$ $=\text{LSTM}(\mathbf{Concat},\mathbf{h}_{t-1},\mathbf{c}_{t-1})$ (3)
$\mathbf{\hat{Y}}$ $=\text{Dense}(\mathbf{h}_{T}),$

### v2

@@ Abstract
The prevalence of hearing aids is increasing. However, optimizing their amplification remains challenging due to the complexity of integrating multiple components in traditional methods. To address this, we present NeuroAMP, a novel deep neural network for end-to-end, personalized amplification in hearing aids. NeuroAMP leverages spectral features and the listener’s audiogram as inputs, and we explore four architectures: Convolutional Neural Network (CNN), Long Short-Term Memory (LSTM), Convolutional Recurrent Neural Network (CRNN), and Transformer. We also introduce Denoising NeuroAMP, an extension that integrates noise reduction with amplification for improved real-world performance. To enhance generalization, we employed a comprehensive data augmentation strategy during training on diverse speech (TIMIT, TMHINT) and music (Cadenza Challenge MUSIC) datasets. Evaluation using the Hearing Aid Speech Perception Index (HASPI), Hearing Aid Speech Quality Index (HASQI), and Hearing Aid Audio Quality Index (HAAQI) shows that the Transformer-based NeuroAMP achieves the best performance, with SRCC scores of 0.9927 (HASQI) and 0.9905 (HASPI) on TIMIT, and 0.9738 (HAAQI) on Cadenza dataset. Notably, the augmentation strategy maintains robust performance on unseen datasets (e.g., VoiceBank-DEMAND, MUSDB18-HQ). Furthermore, Denoising NeuroAMP outperforms both the conventional NAL-R+WDRC method and a two-stage baseline on the VoiceBank-DEMAND dataset, achieving HASPI of 0.90 and HASQI of 0.59. These results highlight the strong potential of NeuroAMP and Denoising NeuroAMP to provide a novel and effective framework for personalized hearing aid amplification.
@@
This study presents NeuroAMP, a novel deep learning-based hearing aid amplifier that replaces complex multi-stage processing with streamlined, end-to-end personalized amplification. We also introduce Denoising NeuroAMP, which integrates noise reduction into the amplification process. Validated on diverse datasets, these models outperform conventional and two-stage methods on standardized metrics such as HASPI and HASQI. Moreover, we believe NeuroAMP’s data-driven approach can advance hearing aid technology by enhancing user-centered assistance and supporting more effective communication for individuals with hearing loss. Future work will explore real-time implementation and incorporate user feedback to further refine personalization.
@@
hearing aids, deep neural network, end-to-end amplification, hearing loss compensation, NeuroAMP, Denoising NeuroAMP
@@ 1 Introduction
Hearing loss, a pervasive global health issue, continues to impact the lives of hundreds of millions, leading to social isolation [1], depression [2], diminished quality of life [3], and even cognitive decline [4]. The World Health Organization (WHO) highlights that approximately 430 million of people worldwide currently face problems related to hearing loss [5, 6]. Despite the availability of hearing aids, which can mitigate these adverse effects, only 16% of individuals with hearing loss in the United States consistently use them [7, 8]. This low adoption rate suggests a notable gap between the potential of hearing aids and their real-world utilization, emphasizing the importance of developing more effective and user-friendly solutions.
@@
Non-use of hearing aids by people with hearing loss can be attributed to a variety of factors, including hearing aids being ineffective in noisy environments, suboptimal sound quality, insufficient benefit, and incompatibility with the individual’s specific type of hearing loss [9]. To address these challenges, various amplifier strategies for hearing aids have been developed [10, 11, 12, 13, 14]. These strategies primarily involve two key components: a prescription fitting formula, which determines the desired insertion gain from the individual’s audiogram, and compression techniques, which are subsequently applied to manage the dynamic range of sounds, as shown in Fig.1. The well-known prescription fitting formulas include NAL-NL1[12], NAL-NL2 [15], NAL-R [10], and DSL m[i/o] [16].
@@
A common problem for people with hearing loss is a reduction in the auditory dynamic range, known as loudness recruitment. In this case, weak sounds may become inaudible, while loud sounds can be uncomfortably intense. Non-linear prescription formulas like NAL-NL1, NAL-NL2, and DSL m[i/o] play a crucial role in addressing the reduced auditory dynamic range through compression. These advanced formulas, however, are primarily used in commercial hearing aids and require licensing, which may limit their broader accessibility. In this study, we use the widely adopted NAL-R prescription formula and incorporate it with Wide Dynamic Range Compression (WDRC) [14] to introduce non-linearity, mimicking the behavior of more advanced systems like NAL-NL1. This approach aligns with the established practice in the open-source OpenMHA platform [17].
@@
Although conventional amplifiers demonstrate notable performance, they often face challenges in adapting to dynamic acoustic environments. These limitations in adaptability and optimization have led to the exploration of alternative approaches. One such promising approach is deep learning, which offers the potential to learn complex patterns from large datasets and create more adaptable systems. Recently, with the availability of huge amounts of training data and their corresponding labels, deep learning models have achieved notable performance in several speech processing applications. Additionally, the integration of deep learning models for speech processing in hearing aid devices has demonstrated robust performance, for example, in speech enhancement [18, 19] and speech assessment tasks [20, 21, 22].
Figure 1: Overview of a typical hearing aid system, illustrating the modular components involved in signal processing, including microphone input, insertion gain, compression, and speaker output.
@@
In this paper, we propose a deep learning-based neural amplifier model called NeuroAMP, designed to provide personalized amplification in an end-to-end manner. Our initial goal is to demonstrate that NeuroAMP can achieve comparable performance to established methods, while also providing the potential for future improvements through its inherent adaptability. Unlike conventional amplifiers (as indicated by the red dashed box in Fig. 1) that combine prescription fitting formulas and compression techniques to amplify audio inputs, NeuroAMP uses a neural network to perform amplification in an end-to-end manner. We hypothesized that the non-linearity of the neural network would allow NeuroAMP to capture a greater variety of acoustic information and more effectively adjust gain and compression based on the user’s specific degree of hearing loss at different frequencies. The development of NeuroAMP involved exploring various neural network architectures, including Convolutional Neural Networks (CNN), Long Short-Term Memory (LSTM), Convolutional Recurrent Neural Networks (CRNN), and Transformer. The training objective of NeuroAMP is to minimize the loss between the estimated amplified audio and the corresponding ground-truth (audio signal processed by conventional insertion gain and compression). To ensure robust performance across diverse conditions, we also apply a data-augmentation strategy using datasets containing speech from multiple languages as well as music. Experimental results confirm the superiority of the Transformer model over other models, achieving the best performance in almost all metrics (with Spearman’s Rank Correlation Coefficient (SRCC) scores of 0.9927 for Hearing Aid Speech Quality Index (HASQI) and 0.9905 for Hearing Aid Speech Perception Index (HASPI) on the TIMIT dataset [23], and 0.9738 for Hearing Aid Audio Quality Index (HAAQI) on the Cadenza challenge dataset [24]). Additionally, we also verify the effectiveness of data augmentation in maintaining good performance on unseen datasets (with SRCC scores of 0.9810 for HASQI and 0.9914 for HASPI on the VoiceBank-DEMAND dataset [25], and 0.9892 for HAAQI on the MUSDB18-HQ dataset [26]). Furthermore, we demonstrate that NeuroAMP is able to mimic conventional amplification methods and provide personalized amplification in an end-to-end manner.
[…]
@@
The remainder of this paper is organized as follows. First, Section 2 reviews related work. Second, Section 3 introduces the proposed methods. Then, Section 4 describes the experimental setup, reports the experimental results, and discusses our findings. Finally, Section 5 concludes this work.
@@ 2.1 Prescription Fitting Formulas
Prescription fitting formulas aim to amplify sound based on the user’s patterns of hearing loss, making speech signals audible across different frequencies while maintaining comfort. These formulas can be broadly categorized into linear and non-linear approaches. Linear amplification, exemplified by methods like NAL-R and NAL-RP [27], provides a constant gain regardless of input volume and is generally preferred for individuals with flatter hearing loss profiles, larger dynamic ranges, and less variability in dynamic range across frequencies. It is also well-suited for those with less varied lifestyles and environments [13]. NAL-R, developed by the Australian National Acoustic Laboratories, is commonly used for mild to moderate hearing loss, while NAL-RP is tailored for severe hearing loss. In contrast, non-linear amplification, such as NAL-NL1, NAL-NL2, and DSL m[i/o], adjusts the gain dynamically based on the input volume, frequency, and the user’s specific hearing loss characteristics. This makes it more suitable for individuals with sloping hearing loss, reduced dynamic ranges, significant variations in dynamic range across frequencies, and more diverse lifestyles and listening environments. NAL-NL1 aims to equalize and normalize the relative volume levels of individual frequencies while optimizing the speech intelligibility for a given volume. NAL-NL2, an improved version of NAL-NL1, considers additional factors in calculating gains, such as gender, age, and whether the native language is tonal, which may affect output volume. NAL-NL2 is also the most widely used fitting formula at present. In contrast, DSL m[i/o] is designed to prevent uncomfortable loudness while using hearing aids and to maximize the audibility of essential messages in conversations. Furthermore, these non-linear formulas utilize compression techniques to manage the reduced auditory dynamic range commonly experienced by individuals with sensorineural hearing loss.
@@ 2.2 Hearing Aid Compression
People with sensorineural hearing loss tend to have a reduced auditory dynamic range. For these individuals, soft sounds are inaudible, while intense sounds are perceived as loudly as they are by a normal-hearing ear. To optimize the use of the residual dynamic range and restore normal loudness perception, most modern hearing aids adopt compression. WDRC is a commonly adopted compression strategy designed to prevent sounds from being either too loud or too soft. The use of WDRC in hearing aids provides several benefits, such as ensuring consistent audibility of speech signals at safe and comfortable hearing levels [14]. Moreover, WDRC involves multiple steps of audio processing, which are [28]:
@@ 3.1 NeuroAMP Architecture
The processing workflow of NeuroAMP involves taking an audio signal and the listener’s audiogram as inputs, processing them through a neural network, and producing an amplified output signal. The detailed architecture of the NeuroAMP model is illustrated in Fig. 2. Specifically, given the input signal $y$, the STFT is applied to extract its time-frequency spectral features ($\textbf{Y}=[\textbf{y}_{1},\dots,\textbf{y}_{t},\dots,\textbf{y}_{T}]$). Simultaneously, the audiogram $z$ is transformed into a vector representation via a dense layer and then replicated into a sequence of length $T$ ($\textbf{Z}=[\textbf{z}_{1},\dots,\textbf{z}_{t},\dots,\textbf{z}_{T}]$). The two input streams are concatenated along the feature dimension and then passed to the subsequent layers of the model. In our setup, we concatenate these two sets of features in the latent space to leverage the neural network’s ability to capture semantic similarities and perform effective feature fusion. This approach has been shown to be effective for building personalized deep-learning models [29]. The feedforward process of NeuroAMP is defined as follows:
[…]
@@
In this work, for the core module in NeuroAMP, we select four neural network architectures, including CNN, LSTM, CRNN, and Transformer.
@@ 3.1.1 Convolutional Neural Network (CNN)
Due to their ability to extract effective features from audio spectrograms [30, 31], CNNs are particularly adept at processing complex audio signals. In NeuroAMP, CNN is used to process spectral features extracted by STFT, with the goal of capturing detailed representations of the frequency components of the audio signal. The operation of CNN in the NeuroAMP model can be described by the following equations:
@@
$\hat{\mathbf{Y}}$ $=\text{Dense}(\text{Flatten}(\mathbf{h}_{L}))$
where $\mathbf{W}_{l}$ and $\mathbf{b}_{l}$ are the convolutional weights and biases of the $l$-th layer, respectively, and $L$ is the number of convolutional layers. This architecture ensures a comprehensive feature extraction process, allowing NeuroAMP to characterize auditory elements more effectively.
@@ 3.1.2 Long Short-Term Memory (LSTM)
In the context of hearing aids, preserving the temporal structure of the audio signal is essential to maintaining speech intelligibility and naturalness. The LSTM is effective at capturing long-range dependencies and temporal patterns [32, 33, 34], which is crucial for processing continuous audio signals and adapting to varying acoustic environments. The process of the LSTM in NeuroAMP is described as follows:
$\mathbf{h}_{t},\mathbf{c}_{t}$ $=\text{LSTM}(\mathbf{Concat}_{t},\mathbf{h}_{t-1},\mathbf{c}_{t-1})$ (3)
$\hat{\mathbf{Y}}$ $=\text{Dense}(\mathbf{h}_{T}),$